OpenStudy (chunk1101):

How do I add parentheses to this equation to make it 19? 5+4*2+6-2*2-1. The answer to the equation is 14 if that helps.

OpenStudy (sshayer):

(5+4)2+6-2*2-1=9*2+6-4-1=18+6-5=?